# Dynamical Dark Energy from Lattice Quantum Gravity

### v2

@@ I Introduction
Although there is no obvious barrier to taking the continuum limit in EDT with the appropriate tuning (other than the practical issues confronting a numerical simulation), more work is needed. To this end, Ref. [19] introduced a new rejection free algorithm that leads to a significant speedup of the simulations in the region of the phase diagram where finer lattices are expected. Ref. [19] also provided a first look at the new, finer ensembles that became feasible to generate using the new algorithm. In this work we turn to a detailed study of these new ensembles. We find that some of our previous methods for determining the lattice spacing within EDT lead to ambiguous results on our new ensembles, which makes these methods insufficient for meeting our target precision. In previous work [15, 21], a diffusion process on the geometries was used to set the relative lattice spacing and to determine the ratio of the dual and direct lattice spacings. Despite the fact that the direct and dual spacings are short-distance quantities, their ratio is fixed by the long-distance emergent properties of the geometry. We show that this ratio can be computed in multiple ways and that the good agreement between these determinations gives confidence both in the results and in the assumptions that go into the determination.
@@
The sign of the effect seen here ensures that it does not violate the null-energy condition. There are strong constraints on theories that are inconsistent with this condition [32, 33, 34], so the fact that our lattice result does not run afoul of it is a significant point in its favor. Although our results are found on lattices that are no larger than several Planck volumes, it is interesting to ask what would happen if the model singled out by this work was assumed to hold throughout cosmic history. One consequence is that the running of the vacuum energy would lead to a large vacuum energy density in the early universe compared to the current era. It turns out that the computed size of the coupling of the model ensures that the effects of the vacuum on the evolution of the universe remain subdominant throughout cosmic history, thus making it compatible with present observational bounds. Nonetheless, the corrections to standard cosmology would not be negligible, with the effects of the running vacuum giving rise to deviations from the $\Lambda$CDM model at the part-per-mil level across many cosmological observables, including the dark energy equation of state.
This paper is organized as follows. In Section II we review the EDT formulation and discuss the details of the simulations. In Section III we introduce two methods for determining the ratio of the direct to dual lattice spacing using the classical de Sitter solution. Details and results from our numerical analysis of this ratio are presented. We also use the semiclassical approximation, where the de Sitter instanton dominates the Euclidean partition function, to extract Newton’s constant at each of our four lattice spacings, thus fixing the lattice scale in terms of the Planck scale. In Section IV we show that the emergent behavior of the EDT geometries is consistent with a cosmological constant that exhibits a power law running with scale. We use our lattice calculation to constrain this behavior, to provide a number of non-trivial cross-checks, and to determine the parameters of the running. Subsection IV.1 provides an introduction to the running vacuum model that we consider, while subsections IV.2, IV.3, IV.4, and IV.5 give an in depth treatment of the numerical analysis used to constrain the parameters of this model. Subsection IV.6 summarizes the model and provides the final parameter values that we obtain from our calculation. In Section V we discuss the implications of the running vacuum model picked out by EDT for observational cosmology. We conclude in Section VI.
@@ II.4 Lattice ensembles
The tuning of the bare lattice parameters described in the previous subsection requires that we simulate close to the phase transition, which means that the cost of generating ensembles grows rapidly with increasing volume. One simple approach to speeding up the lattice generation, which we adopt, is to use multiple streams. We branch multiple streams off of a single thermalized run and then allow the separate streams to run for a full autocorrelation time before we begin to take measurements on them. As noted earlier we estimate the error using a jackknife resampling procedure, and to account for autocorrelation errors we block our data sets, increasing the block size until the jackknife error no longer increases. The integrated autocorrelation time $\tau_{\rm int}$ can then be estimated using
@@
Table 1: Ensembles used in this work. The first column is the $\kappa_{2}$ value of a given lattice spacing. The second is the tuned $\beta$ value of the ensemble. The third is the target volume in four-simplices of the ensemble. The fourth column is the number of configurations generated, and the fifth column is the autocorrelation time on that ensemble as determined by the peak height of the shelling function. The sixth column is the ratio of the direct to dual lattice spacing at the given $\kappa_{2}$ value, and the seventh is the (dual) relative lattice spacing at the given $\kappa_{2}$ value.
@@
Table 1 shows the bare parameters $\kappa_{2}$ and $\beta$ used in the simulations, as well as the target number of four-simplices, $N_{4}$. Once $N_{4}$ is chosen, $\kappa_{4}$ is tuned so that the four-volume fluctuates about the target volume. Table 1 also gives the number of saved configurations in each ensemble, as well as the autocorrelation time $\tau$ of an ensemble, as measured by the Monte Carlo time series for the peak height of the shelling function. The ratio of the direct to dual lattice spacings $a_{\rm lat}/\ell$ and the relative (dual) lattice spacings $\ell_{\rm rel}$ are also given, including the errors in their determinations, for future reference. At our two finest lattice spacings for larger volumes there is some ambiguity in determining the tuned $\beta$ value, so more than one ensemble is used to estimate a tuning error. We discuss in Sections III and IV how we propagate this tuning error to our analyses.
@@
This section presents the method and numerical analysis used to determine the lattice spacing in our calculation. The large number of new ensembles at finer lattices used in this work, as well as our current target precision, makes it necessary to revisit the methods of determining relative and absolute lattice spacings used in previous work. Given that our lattices are not hypercubic, and the geometry at long distance scales is emergent, it is necessary to determine the ratio of direct and dual lattice spacings, $a_{\rm lat}/\ell$ (where $\ell$ is the dual lattice spacing), from the simulations. Some quantities are more naturally measured in terms of one or the other, requiring that we know their relative sizes at each nominal lattice spacing (i.e., at fixed $\kappa_{2}$). We also require the relative lattice spacing $\ell_{\rm rel}$ across different $\kappa_{2}$ values. In previous work [15, 21] we used the return probability of a diffusion process to determine sets of $a_{\rm lat}/\ell$ and $\ell_{\rm rel}$ values. However, we find that there are large finite-volume and discretization effects in the return probability, making it difficult to reach a useful precision with this quantity. In order to improve on this situation, we introduce two new methods for determining $a_{\rm lat}/\ell$ based on the identification of our lattice geometries with de Sitter space. The good agreement between these determinations provides a strong foundation for using our values of $a_{\rm lat}/\ell$ as inputs to determinations of other quantities.
@@ III.1 The classical de Sitter solution
$N_{4}^{\rm shell}=\frac{3}{4}N_{4}\frac{1}{s_{0}N_{4}^{\frac{1}{4}}}\sin^{3}\left(\frac{\tau}{s_{0}N_{4}^{\frac{1}{4}}}\right),$ (13)
where $s_{0}$ is a free parameter and $\tau$ is the Euclidean time in lattice units. It is normalized so that a sum over all time slices gives the total number of four-simplices of the geometry. This expression can be interpreted as Euclidean de Sitter space. It differs from Eq. (12) by a shift in origin of the Euclidean time $\tau$, and by the fact that the shelling function is a three-volume and therefore corresponds to the scale factor cubed. This expression is a fair description of the lattice result for the shelling function, but the presence of discretization effects leads to difficulties in fitting the data, given its precision. In order to obtain good fits, we need to add additional parameters to the expression,
$N_{4}^{\rm shell}=\frac{3}{4}\eta N_{4}\frac{1}{s_{0}N_{4}^{\frac{1}{4}}}\sin^{3}\left(\frac{\tau}{s_{0}N_{4}^{\frac{1}{4}}}+b\right),$ (14)
@@
$N_{4}^{\rm shell}=A_{s}\sin^{3}(B_{s}\tau+C_{s}),$ (15)
where $A_{s}$, $B_{s}$, and $C_{s}$ are free parameters. Because the Euclidean time in the shelling function is measured in dual lattice units, and the lattice scale factor is a measure of the three volume obtained by counting four-simplices in concentric shells, they are not determined in the same units. We can use the fact that we must make contact with a solution to the classical Einstein equations in the appropriate limit in order to find a conversion factor between these two sets of lattice units. If the correct solution is Euclidean de Sitter space, then Eq. 12 provides a relationship between the constant multiplying the trigonometric function and the coefficient in the argument of that function. Then the renormalized units $\ell$ and $\ell_{V}$ must satisfy
@@
where $\ell$ is the dual lattice spacing and $\ell_{V}$ is the unit length derived from taking the cube root of the normalization of the shelling function. These units of measurement can be related via a comparison of the total four-volume of a thin slice of de Sitter space measured either by summing the volumes of four-simplices making up that slice, or by taking the integral over a four-dimensional shell. This leads to the relation
$N_{\rm slice}C_{4}a_{\rm lat}^{4}=2\pi^{2}N_{\rm slice}\ell_{V}^{3}\ell,$ (17)
where $C_{4}=\sqrt{5}/96$, and $N_{\rm slice}$ is the number of four-simplices in a given shell. This expression can be solved for the ratio of direct and dual lattice spacings
@@
Note that $a_{\rm lat}/\ell$ is a renormalized, emergent quantity and is not necessarily given by the expectation from short-distance simplicial geometry. The short-distance geometrical result can be obtained by considering two equilateral $d$-simplexes with edge length $a_{\mathrm{lat}}$ glued together. The distance between the centers is $\sqrt{\frac{2}{d(d+1)}}a_{\mathrm{lat}}$, so that the ratio $a_{\rm lat}/\ell$ is a fixed number $\sqrt{\frac{d(d+1)}{2}}$. However, the effective long-distance value of this ratio is modified from what one would infer from the local geometry, and its value must be determined from the simulations at each lattice spacing.
In summary, $a_{\rm lat}/\ell$ is determined in Method 1 using the fact that the de Sitter radius is measured in two different ways when we construct the shelling function, each in different units. Setting these measures of the radius equal gives a value for the conversion factor $a_{\rm lat}/\ell$. Even though $a_{\rm lat}$ and $\ell$ are short-distance quantities, their ratio is fixed by the long-distance emergent properties of the geometry. Thus, we obtain $a_{\rm lat}/\ell$ using Eqs. (16) and (18) and fits of lattice data to the de Sitter solution.
@@ III.2.1 Determining the Newton constant
Once we have determined $k$ we can obtain $G/a^{2}$, since by Eq. (23), they are inversely proportional. We review here the derivation of that proportionality constant. The same saddle-point expansion used in Eq. (III.2.1) gives for the partition function
@@
with $s_{G}$ the slope determined by a fit to $\kappa_{4}$ as a function of $1/\sqrt{N_{4}}$. A comparison of $G/a_{\rm lat}^{2}$ at different lattice spacings allows us to determine the relative (direct) lattice spacing $a_{\rm rel}$ with reference to a fiducial lattice spacing. The ratios $a_{\rm lat}/\ell$ allow us to compute as well the relative dual lattice spacing $\ell_{\rm rel}$.
@@ III.2.2 Determination of $a_{\rm lat}/\ell$ from Method 2
We can turn this into a relation between the renormalized cosmological constant in dual lattice units and the ratio $a_{\rm lat}/\ell$ by multiplying through by $\ell^{2}$,
@@ III.3.2 $a_{\rm lat}/\ell$ from Method 2
Our second method of determining $a_{\rm lat}/\ell$ from lattice data makes use of Eq. (34), which itself makes use of the relation between the number of four-simplices and the effective cosmological constant of the lattice “universe”. This requires as input $s_{\Lambda}$, which can be obtained from Eq. (33), as well as $\eta$, which can be determined via Eq. (31) from the results of our fits to the classical de Sitter solution. The determination of $s_{\Lambda}$ requires as input the renormalized cosmological constant in dual lattice units, $\Lambda\ell^{2}$, which can be obtained as follows. We relate the lattice shelling function to the scale factor $a$ and its time derivatives. Given this, we can infer the value of the emergent cosmological constant at long distances by constructing the Einstein equation for the scale factor from our lattice data. In Euclidean space, the homogeneous, isotropic Einstein equations in the presence of a positive cosmological constant reduce to
@@
If we identify $A\equiv(N_{4}^{\rm shell})^{1/3}$ as the scale factor determined from the lattice, then we have $a(\tau)=\frac{\ell_{V}}{\ell}A(\tau)$, where $a(\tau)$ is measured in simplex units. To emphasize the dependence on the lattice conversion factors, Eq. (35) written in terms of $A$ becomes
$-\frac{\dot{A}^{2}}{A^{2}}+\frac{\ell^{2}}{\ell_{V}^{2}}\frac{1}{A^{2}}=\frac{\Lambda\ell^{2}}{3},$ (36)
where the ratio $\ell_{V}/\ell$ converts the scale factor $A$ into dual lattice units, with $\ell_{V}/\ell$ related to the ratio of direct and dual lattice spacings $a_{\rm lat}/\ell$ by Eq. 18. The Euclidean time in the time derivative of the first term of Eq. 36 is already in dual lattice units, so the linear combination of terms on the left side of Eq. 36 gives the cosmological constant in dual lattice units $\ell$.

### v3

@@ I Introduction
Although there is no obvious barrier to taking the continuum limit in EDT with the appropriate tuning (other than the practical issues confronting a numerical simulation), more work is needed. To this end, Ref. [19] introduced a new rejection free algorithm that leads to a significant speedup of the simulations in the region of the phase diagram where finer lattices are expected. Ref. [19] also provided a first look at the new, finer ensembles that became feasible to generate using the new algorithm. In this work we turn to a detailed study of these new ensembles. We find that some of our previous methods for determining the lattice spacing within EDT lead to ambiguous results on our new ensembles, which makes these methods insufficient for meeting our target precision. In previous work [15, 21], a diffusion process on the geometries was used to set various relative lattice spacings. We revisit these determinations here, especially the ratio between the bare direct lattice spacing and the renormalized dual lattice spacing. We show that this ratio can be computed in multiple ways and that the good agreement between these determinations gives confidence both in the results and in the assumptions that go into the determination.
@@
The sign of the effect seen here ensures that it does not violate the null-energy condition. There are strong constraints on theories that are inconsistent with this condition [32, 33, 34], so the fact that our lattice result does not run afoul of it is a significant point in its favor. Although our results are found on lattices that are no larger than several Planck volumes, it is interesting to ask what would happen if the model singled out by this work was assumed to hold throughout cosmic history. One consequence is that the running of the vacuum energy would lead to a large vacuum energy density in the early universe compared to the current era. It turns out that the computed size of the coupling of the model ensures that the effect of the new vacuum dynamics on the evolution of the universe remains subdominant throughout cosmic history, thus making it compatible with present observational bounds. Nonetheless, the corrections to standard cosmology would not be negligible, with the effects of the running vacuum giving rise to deviations from the $\Lambda$CDM model at the part-per-mil level across many cosmological observables, including the dark energy equation of state.
This paper is organized as follows. In Section II we review the EDT formulation and discuss the details of the simulations. In Section III we introduce two methods for determining the ratio of the direct to (renormalized) dual lattice spacing using the classical de Sitter solution. Details and results from our numerical analysis of this ratio are presented. We also use the semiclassical approximation, where the de Sitter instanton dominates the Euclidean partition function, to extract Newton’s constant at each of our four lattice spacings, thus fixing the lattice scale in terms of the Planck scale. In Section IV we show that the emergent behavior of the EDT geometries is consistent with a cosmological constant that exhibits a power law running with scale. We use our lattice calculation to constrain this behavior, to provide a number of non-trivial cross-checks, and to determine the parameters of the running. Subsection IV.1 provides an introduction to the running vacuum model that we consider, while subsections IV.2, IV.3, IV.4, and IV.5 give an in depth treatment of the numerical analysis used to constrain the parameters of this model. Subsection IV.6 summarizes the model and provides the final parameter values that we obtain from our calculation. In Section V we discuss the implications of the running vacuum model picked out by EDT for observational cosmology. We conclude in Section VI.
@@ II.4 Lattice ensembles
The tuning of the bare lattice parameters described in the previous subsection requires that we simulate close to the phase transition, which means that the cost of generating ensembles grows rapidly with increasing volume. One simple approach to speeding up the lattice generation, which we adopt, is to use multiple streams. We branch multiple streams off of a single thermalized run and then allow the separate streams to run for a full autocorrelation time before we begin to take measurements on them. As noted in the previous subsection we estimate the error using a jackknife resampling procedure, and to account for autocorrelation errors we block our data sets, increasing the block size until the jackknife error no longer increases. The integrated autocorrelation time $\tau_{\rm int}$ can then be estimated using
@@
Table 1: Ensembles used in this work. The first column is the $\kappa_{2}$ value of a given lattice spacing. The second is the tuned $\beta$ value of the ensemble. The third is the target volume in four-simplices of the ensemble. The fourth column is the number of configurations generated, and the fifth column is the autocorrelation time on that ensemble as determined by the peak height of the shelling function. The sixth column is the ratio of the direct to renormalized dual lattice spacing at the given $\kappa_{2}$ value, and the seventh is the (renormalized dual) relative lattice spacing at the given $\kappa_{2}$ value.
@@
Table 1 shows the bare parameters $\kappa_{2}$ and $\beta$ used in the simulations, as well as the target number of four-simplices, $N_{4}$. Once $N_{4}$ is chosen, $\kappa_{4}$ is tuned so that the four-volume fluctuates about the target volume. Table 1 also gives the number of saved configurations in each ensemble, as well as the autocorrelation time $\tau$ of an ensemble, as measured by the Monte Carlo time series for the peak height of the shelling function. The ratio of the direct to renormalized dual lattice spacings $a_{\rm lat}/\ell$ and the relative (renormalized dual) lattice spacings $\ell_{\rm rel}$ are also given, including the errors in their determinations, for future reference. At our two finest lattice spacings for larger volumes there is some ambiguity in determining the tuned $\beta$ value, so more than one ensemble is used to estimate a tuning error. We discuss in Sections III and IV how we propagate this tuning error to our analyses.
@@
This section presents the method and numerical analysis used to determine the lattice spacing in our calculation. The large number of new ensembles at finer lattices used in this work, as well as our current target precision, makes it necessary to revisit the methods of determining relative and absolute lattice spacings used in previous work. Given that our lattices are not hypercubic, and the geometry at long distance scales is emergent, it is necessary to determine the ratio, $a_{\rm lat}/\ell$, where $\ell$ is a distance scale that we identify as a renormalized dual lattice spacing, from the simulations. Some quantities are more naturally measured in terms of $a_{\rm lat}$ or in terms of $\ell$, requiring that we know their relative sizes at each nominal lattice spacing (i.e., at fixed $\kappa_{2}$). We also require the relative lattice spacing $\ell_{\rm rel}$ across different $\kappa_{2}$ values. In previous work [15, 21] we used the return probability of a diffusion process to determine sets of $a_{\rm lat}/\ell$ and $\ell_{\rm rel}$ values. However, we find that there are large finite-volume and discretization effects in the return probability, making it difficult to reach a useful precision with this quantity. In order to improve on this situation, we introduce two new methods for determining $a_{\rm lat}/\ell$ based on the identification of our lattice geometries with de Sitter space. The good agreement between these determinations provides a strong foundation for using our values of $a_{\rm lat}/\ell$ as inputs to determinations of other quantities.
@@ III.1 The classical de Sitter solution
$N_{4}^{\rm shell}=\frac{3}{4}N_{4}\frac{1}{s_{0}N_{4}^{\frac{1}{4}}}\sin^{3}\left(\frac{\tau/\ell}{s_{0}N_{4}^{\frac{1}{4}}}\right),$ (13)
where $s_{0}$ is a free parameter and $\tau/\ell$ is the Euclidean time in renormalized dual lattice ($\ell$) units. $N_{4}^{\rm shell}$ is normalized so that a sum over all time slices gives the total number of four-simplices of the geometry. This expression can be interpreted as Euclidean de Sitter space. It differs from Eq. (12) by a shift in origin of the Euclidean time $\tau$, and by the fact that the shelling function is a three-volume and therefore corresponds to the scale factor cubed. This expression is a fair description of the lattice result for the shelling function, but the presence of discretization effects leads to difficulties in fitting the data, given its precision. In order to obtain good fits, we need to add additional parameters to the expression,
$N_{4}^{\rm shell}=\frac{3}{4}\eta N_{4}\frac{1}{s_{0}N_{4}^{\frac{1}{4}}}\sin^{3}\left(\frac{\tau/\ell}{s_{0}N_{4}^{\frac{1}{4}}}+b\right),$ (14)
@@
$N_{4}^{\rm shell}=A_{s}\sin^{3}(B_{s}\tau/\ell+C_{s}),$ (15)
where $A_{s}$, $B_{s}$, and $C_{s}$ are free parameters. Because the Euclidean time in the shelling function is measured in dual lattice units, and the lattice scale factor is a measure of the three volume obtained by counting four-simplices in concentric shells, they are not determined in the same units. We can use the fact that we must make contact with a solution to the classical Einstein equations in the appropriate limit in order to find a conversion factor between these two sets of lattice units. If the correct solution is Euclidean de Sitter space, then Eq. 12 provides a relationship between the constant multiplying the trigonometric function and the coefficient in the argument of that function. Then the distance units $\ell$ and $\ell_{V}$ must satisfy
@@
where $\ell$ is the renormalized dual lattice spacing and $\ell_{V}$ is the unit length derived from taking the cube root of the normalization of the shelling function. These units of measurement can be related via a comparison of the total four-volume of a thin slice of de Sitter space measured either by summing the volumes of four-simplices making up that slice, or by taking the integral over a four-dimensional shell. This leads to the relation
$N_{4}^{\rm shell}(\tau)C_{4}a_{\rm lat}^{4}=2\pi^{2}N_{4}^{\rm shell}(\tau)\ell_{V}^{3}\ell,$ (17)
where $C_{4}=\sqrt{5}/96$, and $N_{4}^{\rm shell}(\tau)$ is the number of four-simplices in a given shell. Eqs. (16) and (17) may be taken to implicitly define $\ell$ and $\ell_{V}$. Eq. (17) can then be solved for the ratio of direct and renormalized dual lattice spacings
@@
Note that $\ell$, and thus $a_{\rm lat}/\ell$, is an emergent quantity not necessarily given by the expectation from short-distance simplicial geometry. If we denote the short-distance dual lattice spacing by $\tilde{\ell}$, then it can be shown that $\tilde{\ell}$ is related to $a_{\rm lat}$ as follows. Given two equilateral $d$-simplexes with edge length $a_{\mathrm{lat}}$ glued together, the distance between the centers is $\sqrt{\frac{2}{d(d+1)}}a_{\mathrm{lat}}$, so that the ratio $a_{\rm lat}/\tilde{\ell}$, is a fixed number $\sqrt{\frac{d(d+1)}{2}}$. The effective long-distance $\ell$ differs from $\tilde{\ell}$, and its value must be determined from the simulations at each lattice spacing. With $\ell$ given implicitly by Eqs. (16) and (17), it is defined in an operational way that is not altered by the use of degenerate triangulations.
In summary, $a_{\rm lat}/\ell$ is determined in Method 1 using the fact that the de Sitter radius is measured in two different ways when we construct the shelling function, each in different units. Setting these measures of the radius equal gives a value for the conversion factor $a_{\rm lat}/\ell$. Thus, we obtain $a_{\rm lat}/\ell$ using Eqs. (16) and (18) and fits of lattice data to the de Sitter solution.
@@ III.2.1 Determining the Newton constant
Once we have determined $k$ we can obtain $G/a_{\rm lat}^{2}$, since by Eq. (23), they are inversely proportional. We review here the derivation of that proportionality constant. The same saddle-point expansion used in Eq. (III.2.1) gives for the partition function
@@
with $s_{G}$ the slope determined by a fit to $\kappa_{4}$ as a function of $1/\sqrt{N_{4}}$. A comparison of $G/a_{\rm lat}^{2}$ at different lattice spacings allows us to determine the relative (direct) lattice spacing $a_{\rm rel}$ with reference to a fiducial lattice spacing. The ratios $a_{\rm lat}/\ell$ allow us to compute as well the relative renormalized dual lattice spacing $\ell_{\rm rel}$.
@@ III.2.2 Determination of $a_{\rm lat}/\ell$ from Method 2
We can turn this into a relation between the renormalized cosmological constant in $\ell$ units and the ratio $a_{\rm lat}/\ell$ by multiplying through by $\ell^{2}$,
@@ III.3.2 $a_{\rm lat}/\ell$ from Method 2
Our second method of determining $a_{\rm lat}/\ell$ from lattice data makes use of Eq. (34), which itself makes use of the relation between the number of four-simplices and the effective cosmological constant of the lattice “universe”. This requires as input $s_{\Lambda}$, which can be obtained from Eq. (33), as well as $\eta$, which can be determined via Eq. (31) from the results of our fits to the classical de Sitter solution. The determination of $s_{\Lambda}$ requires as input the renormalized cosmological constant in $\ell$ units, $\Lambda\ell^{2}$, which can be obtained as follows. We relate the lattice shelling function to the scale factor $a$ and its time derivatives. Given this, we can infer the value of the emergent cosmological constant at long distances by constructing the Einstein equation for the scale factor from our lattice data. In Euclidean space, the homogeneous, isotropic Einstein equations in the presence of a positive cosmological constant reduce to
@@
The scale factor measured in $\ell$ units is obtained directly from the lattice scale factor
$\frac{a(\tau)}{\ell}=\frac{\ell_{V}}{\ell}\frac{A(\tau)}{\ell_{V}},$ (36)
where we introduce the notation
$A(\tau)/\ell_{V}\equiv(N_{4}^{\rm shell}(\tau))^{\frac{1}{3}},$ (37)
to relate the lattice scale factor to the shelling function, emphasizing the dependence on the lattice spacing conversion factors. Then Eq. (35) can be written in terms of $A$,
$-\frac{\dot{A}^{2}}{A^{2}}+\frac{\ell^{2}}{\ell_{V}^{2}}\frac{\ell_{V}^{2}}{A^{2}}=\frac{\Lambda\ell^{2}}{3},$ (38)
where the ratio $\ell_{V}/\ell$ converts the scale factor $A/\ell_{V}$ into renormalized dual lattice ($\ell$) units, with $\ell_{V}/\ell$ related to the ratio $a_{\rm lat}/\ell$ by
The Euclidean time in the time derivative of the first term of Eq. 38 is already in $\ell$ units, so the linear combination of terms on the left side of Eq. 38 gives the cosmological constant also in $\ell$ units.